OpenStudy (anonymous):

Say that a and b are the lengths of the legs of a right triangle and c is the length of the hypotenuse. Find c if a = 16 and b = 30.

OpenStudy (anonymous):

a. c = 17 c. c = 34 b. c = 1,156 d. c = 68

OpenStudy (solomonzelman):

\[\huge\color{blue}{a^2+b^2=c^2}~~~~~~~c=hypotenuse\]

OpenStudy (anonymous):

so it would be d

OpenStudy (solomonzelman):

Nope, it's not B.

OpenStudy (solomonzelman):

Depends on which one is B in your q. I know it's 34.

OpenStudy (anonymous):

its c 34

OpenStudy (solomonzelman):

Yes, \[16^2+30^2=c^2,~~~~~~~~~1156=c^2,~~~~~~~~~~~\sqrt{1156}=\sqrt{c^2},~~~~~~~c=34\]

OpenStudy (anonymous):

thank you for formula that helps

OpenStudy (solomonzelman):

Anytime!